OpenStudy (anonymous):

elimination using multiplication 2x+5y=3 -x+3y=-7

OpenStudy (anonymous):

2x+5y=3 ( -x+3y=-7 ) * 2 ------------ 2x+5y=3 -2x+6y=-14 (eliminate x) ------------ 11y=-11 therefore y = -1 and 2x+5(-1)=3 -> 2x -8=3 -> 2x=11 -> x =11/2

OpenStudy (anonymous):

Thanks